what is the perimeter of the rectangle if the length is 2.54cm. and the width is 1.9cm.?

Answer by mananth(16949)

perimeter = 2*(l+w)
l=2.54
w=1.9
Perimeter = 2*(2.54+1.9)
=2*(4.44)
= 8.88 cm